OpenStudy (anonymous):

2x + 3y = 6 5x + 2y = 4 Which of the following equations could be the result of multiplication and addition to eliminate a variable in the system of equations?

OpenStudy (danjs):

you can multiply a whole equation by some number and it will still be the same equation how can you get the x term or the y term opposite signs of each other? so when you add the two equations, one variable goes to 0x or 0y

OpenStudy (danjs):

like 2 times the first equation and (-3) times the second equation , try that out

OpenStudy (anonymous):

Oh! Okay. Let me try it.. pls don't leave lol

OpenStudy (danjs):

i was about to, but i will stay.. hah

OpenStudy (anonymous):

How did you think of using -3 for the second one?

OpenStudy (danjs):

type out the two new equations

OpenStudy (anonymous):

Okay

OpenStudy (danjs):

i was looking at the y variables, 3y and 2y can both be turned into 6y , need the negative to get the -6y for the second

OpenStudy (anonymous):

4x + 6y = 12 -15x -6y = -12

OpenStudy (anonymous):

OH! I understand what you mean

OpenStudy (danjs):

looks good, you can also add equations together in systems like this

OpenStudy (anonymous):

But, in this case I'm substracting right?

OpenStudy (anonymous):

I got -11x = 0 by subtracting

OpenStudy (danjs):

4x + 6y = 12 -15x -6y = -12 ----------------- + nah just think of addition all the time, subtraction is the sam, plus a negative

OpenStudy (danjs):

good, so you can get the x value

OpenStudy (danjs):

x = 0/-11 = 0

OpenStudy (anonymous):

It's 0 lol

OpenStudy (anonymous):

Great! Thank you!!

OpenStudy (danjs):

use that in either equation and solve for the y value now

OpenStudy (anonymous):

Thanks :)

OpenStudy (danjs):

the solution is the point (x,y) where thosew 2 lines intersect

OpenStudy (danjs):

both have same x and same y

OpenStudy (danjs):

welcome

OpenStudy (anonymous):

:D

OpenStudy (danjs):

try 5xfirst and -2 x second, the x goes away this time, and you should get the same answer for x and y

OpenStudy (danjs):

just remember the general rules for these systems of lines, 2 or more equations 1 - multiply by a constant 2- add equations together 3 - change order of equations that pretty much it